OpenStudy (anonymous):

help with my answers corrections medal plus fan . Which of the following statements have the same result? Explain each step in solving each one. I f(3) when f(x) = 2x + 2 II f-1(4) when f(x) = 3x minus 4, all over 5 III y + 10 = 2y + 1

OpenStudy (anonymous):

I. f(3) when f(x)= 2x+2 f(x)=2x+2 f(3)=2(3)+2= 8 II. f^-1(4) when f(x)= 3x-4 all over 5 f(x)= 3x-4 all over 5 To find f^-1(x) just swap the variables to solve for x: y= 3x-4 all over 5 swap variables: x= 3y-4 all over 5 solve for x 5x=3y-4 5x+4= 3y y= 5x+4 all over 3 f^-1(x)= 5x+4 all over 3 5*4= 20 20+4= 24 24/3= 8 III. y+10 = 2y+1 y+10= 2y+1 10= y+1 y= 9 I and II equals to 8, which for this question these functions has the same statements, while III equals to 9.

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Gokuporter

ganeshie8 (ganeshie8):

Looks great ! good job !!

OpenStudy (anonymous):

okay im stuck on another 1 please help

ganeshie8 (ganeshie8):

okay

OpenStudy (anonymous):

Emma wants to compare some cell phone plans by graphing their cost per text message. Mobilama charges a flat rate of $30 for unlimited texts. Celly-On-The-Go has no set-up charge, but it does charge $0.10 per text. Ringit Wireless only charges $0.05 per text, but it has an account charge of $15. Explain to Emma how she can graph these cell phone plans and describe what key features of the graphs she should consider when making a decision.

ganeshie8 (ganeshie8):

setup the equations first

ganeshie8 (ganeshie8):

`Mobilama charges a flat rate of $30 for unlimited texts. ` \[y = 30\]

ganeshie8 (ganeshie8):

`Celly-On-The-Go has no set-up charge, but it does charge $0.10 per text. ` \[y = 0.1x\]

ganeshie8 (ganeshie8):

can you setup the equation for third plan ?

OpenStudy (anonymous):

no, i think

ganeshie8 (ganeshie8):

did u get how the above two equations were setup ?

OpenStudy (anonymous):

yehh

ganeshie8 (ganeshie8):

good, see if u can try and setup the third equation same way... give it a try :)

ganeshie8 (ganeshie8):

`Ringit Wireless only charges $0.05 per text, but it has an account charge of $15.`

OpenStudy (anonymous):

kk

ganeshie8 (ganeshie8):

translate that word problem into an equation

OpenStudy (anonymous):

can x be involved

ganeshie8 (ganeshie8):

say, x = number of texts `Ringit Wireless only charges $0.05 per text,` so cost of "x" texts would be : `0.05x`

OpenStudy (anonymous):

ohh i see

ganeshie8 (ganeshie8):

but since they also charge a fixed account charge of $15, the expression for total cost wud be : `0.05x + 15`

OpenStudy (anonymous):

so the varyiabe would either be attached or equal to to the number

OpenStudy (anonymous):

variable*

ganeshie8 (ganeshie8):

Here are the equations for `total cost` of each of 3 plans : Mobilama : \(y=30\) Celly : \(y = 0.1x\) Right Wireless : \(y = 0.05x + 15\)

ganeshie8 (ganeshie8):

btw, the variable `y` represents "total cost" the variable `x` represents "number of text messages"

ganeshie8 (ganeshie8):

graph them in geogebra and analyze them

ganeshie8 (ganeshie8):

you need to look for which graph stays low (that is less cost) and in what intervals (upto how many messages)

ganeshie8 (ganeshie8):

do u have geogebra on ur computer ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

http://www.geogebra.org/webstart/geogebra.html

OpenStudy (anonymous):

this is my moms computer i cant

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

click that link, u can graph it online

ganeshie8 (ganeshie8):

enter the equations in `Input` bar

OpenStudy (anonymous):

i know but it says download and im not alllowed to download anything

OpenStudy (anonymous):

okay @ganeshie8 i got it now what do i do, do i plug in the expressions that you taught me and try to find the lowest line

OpenStudy (anonymous):

im lost on how to work this

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

@whpalmer4 needd help please help

OpenStudy (anonymous):

@blueweek77

ganeshie8 (ganeshie8):

yes ! graph them and get a feel of how the lines look first, and see which line stays low..

OpenStudy (anonymous):

i dont know how cuhz at the bottom i put in y= 30 and it went str8 to the side

ganeshie8 (ganeshie8):

it shows the graph on right side also

OpenStudy (anonymous):

hold up i think i got it but whe im done can you look over my answers for this question

ganeshie8 (ganeshie8):

sure :) u need to press `ctrl -` (control minus)

ganeshie8 (ganeshie8):

to zoom out

ganeshie8 (ganeshie8):

once u graph all the 3 lines, analyze them a bit and attach the snapshot here ok ?

OpenStudy (anonymous):

okay hold up

ganeshie8 (ganeshie8):

take ur time

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Excellent ! now look at the lines, which plan is cheap ?

OpenStudy (anonymous):

y=0.05x+15

ganeshie8 (ganeshie8):

think again

OpenStudy (anonymous):

.01 right i misread it

ganeshie8 (ganeshie8):

whats the full equation ?

OpenStudy (anonymous):

y=0.1x

ganeshie8 (ganeshie8):

and why is it cheap ? explain

ganeshie8 (ganeshie8):

you're right ! could u explain why do u think its the cheapest option ?

OpenStudy (anonymous):

because unlike the others it starts in the middle but falls rapidly i think

ganeshie8 (ganeshie8):

All lines are rising up right ?

ganeshie8 (ganeshie8):

the y axis represents COST, and the x axis represents number of text messages

ganeshie8 (ganeshie8):

As u text more, u need to pay more as well right ?

ganeshie8 (ganeshie8):

Notice that the \(\color{Red}{red}\) line has LOW value when u limit ur texts to `less than 300`

ganeshie8 (ganeshie8):

so thats the cheapest option for you if ur texts are `less than 300`

ganeshie8 (ganeshie8):

okay wid that ?

OpenStudy (anonymous):

yes i jus have to add a little more to my answer then can you correct it

ganeshie8 (ganeshie8):

yes and think about which option is cheap if ur texts are more than 300

ganeshie8 (ganeshie8):

also u may write about which option is costly based on ur number of texts

OpenStudy (anonymous):

Mobilama charges a flat rate of $30 for unlimited texts. y= 30 Celly-On-The-Go has no set-up charge, but it does charge. y= 0.1x Ringit Wireless only charges $0.05 per text. x= 0.05 but since they also charge a fixed account charge of $15, the expression for total cost would be : 0.05x+15 the total costs of each 3 plans would be: Mobilama : y=30 Celly : y=0.1x Right Wireless : y=0.05x+15 Basically when you do word problems x and y is built in the words. So in word y represents the "total cost" and x represents " number of text messages". so on a graph y=0.1x would be the lowest line than the others, even though they all rise up, the more you text, they more they take, because your paying more money, so basically the others has high values and the only lowest value on the graph is y=0.1x.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

y=0.1x is cheap only if ur texts are less than 300

ganeshie8 (ganeshie8):

look at the graph, zoom out and go to the far right when x = 300

ganeshie8 (ganeshie8):

the lines are crossing one another at that point

ganeshie8 (ganeshie8):

y = 0.1x is not longer low after x = 300, right ?

OpenStudy (anonymous):

yehh there both at the same spot

OpenStudy (anonymous):

after x it rises just like the others ohh i see hold up

OpenStudy (anonymous):

so on a graph y=0.1x would be the lowest line than the others, even though they all rise up, the more you text, they more they take, because your paying more money, but at the reaching point of 300 ( on graph ), y=0.1x is no longer the lowest line after x=300.

OpenStudy (anonymous):

so is this the problem @wolf1728

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (wolf1728):

What exactly do you mean "Is this the problem"?

OpenStudy (anonymous):

nvm about that 1 can you help with this 1

OpenStudy (anonymous):

Megan and Julie are stuck simplifying radical expressions. Megan has to simplify the quantity of x to the one third power, over x to the one twelfth power. Julie has to simplify the thirty second root of the quantity of x times x to the second times x to the fifth. Using full sentences describe how to fully simplify Megan and Julie’s expressions. Describe if Megan and Julie started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (anonymous):

pleasee

OpenStudy (wolf1728):

Well I might have to make a graphic to avoid: x^1/3/x^1/12 that sort of thing.

OpenStudy (anonymous):

you can do anyhing you want, :) i just want to the right answer and learn at the same time

OpenStudy (wolf1728):

Okay here's the problem (I think)

OpenStudy (wolf1728):

So, is Julie's calculation supposed to be the 32nd root of x^20 ?